OpenStudy (anonymous):

find the domain of f(x) = ln(x^3 - x)

OpenStudy (anonymous):

hint : the domain of ln(x) is x>0

OpenStudy (anonymous):

So, following the hint, \[x^3-x>0 \implies x(x^2-1)=x(x=1)(x-1)>0 \implies x \in (-\infty,-1) \cup (1,\infty)\]

OpenStudy (anonymous):

Typo in my equation above: Second of three factors is (x+1).

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

No sweat. Do math every day.